OpenStudy (anonymous):

At his normal speed, Shirish can travel 18 km downstream in a fast flowing stream in 9 hours less than what he takes to travel the same distance upstream. The downstream trip would take 1 hour less than what the upstream trip would take provided he doubles his rate of rowing. What is the speed of the stream in km/hr?

OpenStudy (anonymous):

18/(v + s) = 18/(v - s) - 9 18/(2v + s) = 18/(2v - s) - 1 http://www.wolframalpha.com/input/?i=18%2F%28v+%2B+s%29+%3D+18%2F%28v+-+s%29+-+9+and+18%2F%282v+%2B+s%29+%3D+18%2F%282v+-+s%29+-+1 s = 20/3 = 6.67 km/h